# Congestion-Aware Path Re-routing Strategy for Dense Urban Airspace

### v1

@@ VI Conclusion
In this paper, we present a heading-constrained airspace design for the altitude-layered urban airspace and propose a novel preference-based distributed congestion mitigation strategy. The proposed strategy offers a unified framework for handling contingencies such as congestion in the urban airspace, conflicts among exogenous UAS traffic streams, and static obstacle avoidance. The advantages of the proposed approach are many-fold. The airspace design is the responsibility of UTM which could be done offline and the UTM parameters could be communicated to the UAS as needed. The strategy is decentralized with individual UAS being responsible for mitigating congestion and minimal role played by UTM. Hence, the strategy is scalable to the rapidly growing number of UAS. The strategy improves UAS intrinsic safety by guaranteeing no more than $M$ UAS are present in any zone, also reducing the number of UAS conflicts. However, this comes at the expense of a minimal time delay and energy costs introduced in the UAS flight path associated with lateral and diagonal maneuvers. The emerging traffic behavior when every UAS employs the proposed strategy closely resembles a parallel air-corridor network, with the number of active air-corridors (referred to as traffic stream spread) adapting to the UAS demand. The UAS traffic is directional, thus no compromise on the UTM throughput. The paper also presents queueing theory models for estimating the expected traffic spread for any given UAS demand. The estimates help the UTM reserve the airspace beforehand for any foreseen congestion. Extending the above strategy to the general scenario where more than two exogenous UAS traffic streams intersect (may not be perpendicular) and affect each other would be the future avenue of this work. Further, proposing a probabilistic confidence definition for congestion instead of the present deterministic Definition 1 would help further reduce the internal conflicts. There may be other factors that affect the UAS preference to avoid or enter a congested region, that can be incorporated into the above strategy.
@@ Acknowledgment
The research reported here was funded by the Commonwealth Scholarship Commission and the Foreign, Commonwealth and Development Office in the UK. Sajid Ahamed is grateful for their support. All views expressed here are those of the author(s) not the funding body.

### v2

@@ 6 Conclusion
In this paper, we present an airspace design for the altitude-layered urban airspace and propose a novel preference-based distributed congestion mitigation strategy. The proposed strategy offers a unified framework for handling contingencies such as congestion in the urban airspace, conflicts among exogenous UAS traffic streams, and static obstacle avoidance. The advantages of the proposed approach are many-fold. The airspace design is the responsibility of UTM, which could be done offline, and the UTM parameters could be communicated to the UAS as needed. The strategy is decentralized, with individual UAS being responsible for mitigating congestion with minimal role played by UTM. Hence, the strategy is scalable to the rapidly growing number of UAS. The strategy improves UAS intrinsic safety by guaranteeing no more than $M$ UAS are present in any zone, also reducing the number of UAS conflicts. However, this comes at the expense of a minimal time delay and energy costs introduced in the UAS flight path associated with lateral and diagonal maneuvers. The emerging traffic behavior when every UAS employs the proposed strategy closely resembles a parallel air-corridor network, with the number of active air-corridors (referred to as traffic stream spread) adapting to the UAS demand. The UAS traffic is directional; thus, there is no compromise on the UTM throughput. The paper also presents queueing theory models for estimating the expected traffic spread for any given UAS demand. The estimates help the UTM reserve the airspace beforehand for any foreseen congestion. Extending the above strategy to the general scenario where more than two exogenous UAS traffic streams intersect (may not be perpendicular) and affect each other would be the future direction for this work. Further, proposing a probabilistic confidence definition for congestion instead of the present deterministic Definition 1 would help further reduce the internal conflicts. There may be other factors that affect the UAS preference to avoid or enter a congested region that can be incorporated into the above strategy.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ Acknowledgment
The research reported here was funded by the Commonwealth Scholarship Commission and the Foreign, Commonwealth and Development Office in the UK. Sajid Ahamed is grateful for their support. All views expressed here are those of the author(s), not the funding body.
Appendix
Theorem 1.
For Stream$\{0\}$ queuing system, the polynomial equation in $\Theta_{0}$ has at least one real root in $[0,1]$
Proof.
Consider a simplistic case where the queuing system has $L=3$ timeslots queuing deadline. The PGF of the arrival process is given by $A(z)=(1-\lambda)+\lambda z$. Recursively, substituting $V_{j}(z)$ in $V_{j+1}(z)$, $j=1,2,$ we get,
$V_{1}(z)=A(z)\Theta_{0}+(1-\Theta_{0})\ \ ,\ \ V_{2}(z)=A(z)V_{1}(z)\Theta_{0}\ +\left[V_{1}(0)+\left(\cfrac{V_{1}(z)-V_{1}(0)}{z}\right)A(z)\right](1-\Theta_{0})$ (53)
$V_{3}(z)=A(z)V_{2}(z)\Theta_{0}\ +\left[V_{2}(0)+\left(\cfrac{V_{2}(z)-V_{2}(0)}{z}\right)A(z)\right](1-\Theta_{0})$ (54)
The computation of $V_{3}(z)$ requires $V_{2}(0)$, which on substituting $z=0$ in Eqn. 53 is found to be indeterminate ($/0$ 0) because of the $(V_{1}(z)-V_{1}(0))/z$ term. However, when we substitute $A(z)$, $V_{1}(z)$, $V_{1}(0)$ and expand the term,
[…]
we observe that the numerator has a zero at $z=0$. Similarly, in Eqn. 54, the term $V_{2}(z)-V_{2}(0)=z(\ z\lambda^{2}\Theta_{0}^{2}+2\lambda(1-\lambda)\Theta_{0}^{2}+2\lambda\Theta_{0}(1-\Theta_{0})\ )$ has a zero at $z=0$. Irrespective of the number of times $V_{j}(z)$ is recursively substituted in $V_{j+1}(z)$, $j=1,2,...,L-1$, there would be pole-zero cancellations in $(V_{j}(z)-V_{j}(0))/z$.
The pole-zero cancellation is a consequence of Rouché’s theorem. Rouché’s theorem states that if $f(z)$ and $g(z)$ are two analytic functions inside and on a simple closed curve $C$ such that $|f(z)|>|g(z)|$ for all $z\in C$, then both $f(z)$ and $f(z)+g(z)$ have the same number of zeros inside $C$. Consider the simple curve $|z|=1$.
$V_{j}(z)\ =\ A(z)V_{j-1}(z)\Theta_{0}\ +\left[V_{j-1}(0)+\left(\cfrac{V_{j-1}(z)-V_{j-1}(0)}{z}\right)A(z)\right](1-\Theta_{0})$
$=\ V_{j-1}(0)(1-\Theta_{0})\left[1+\cfrac{A(z){{[}}(V_{j-1}(z)-V_{j-1}(0))(1-\Theta_{0})+z\Theta_{0}V_{j-1}(z){{]}}}{zV_{j-1}(0)(1-\Theta_{0})}\right]=\ V_{j-1}(0)(1-\Theta_{0})\left[1+\cfrac{g(z)}{f(z)}\right]$ (56)
When $|z|=1$, we have $|f(z)|=V_{j-1}(0)(1-\Theta_{0})$. If $z=1$, then ${|}A(z){|}=|1-\lambda+\lambda z|=1,\ {|}V_{j-1}(z){|}=1$.
${|}g(z){|}={|}A(z){|}{|}(V_{j-1}(z)-V_{j-1}(0))(1-\Theta_{0})+z\Theta_{0}V_{j-1}(z){|}=1-V_{j-1}(0)(1-\Theta_{0})<V_{j-1}(0)(1-\Theta_{0})$ (57)
If $|z|=1$ but $z\neq 1$, then $|A(z)|<1-\lambda+\lambda|z|<1$, $|V_{j-1}(z)|<1$.
${|}g(z){|}$ $={|}A(z){|}{|}(V_{j-1}(z)-V_{j-1}(0))(1-\Theta_{0})+z\Theta_{0}V_{j-1}(z){|}$
$<{|}A(z){|}\left(\ {|}V_{j-1}(z){|}(1-\Theta_{0})+V_{j-1}(0)(1-\Theta_{0})+\Theta_{0}{|}V_{j-1}(z){|}\ \right)$
$<{|}A(z){|}{|}V_{j-1}(z){|}+{|}A(z){|}V_{j-1}(0)(1-\Theta_{0})\ <\ V_{j-1}(0)(1-\Theta_{0})$ (58)
Since $|f(z)|>|g(z)|$ for all $|z|=1$, the numerator $f(z)+g(z)$ and denominator $f(z)$ have same number of zeros inside the region $|z|<1$. Let $z_{1},z_{2}$ be the zeros of $f(z)$ and $g(z)$, respectively. By the $j^{th}$ recursion, the term $(z_{2}/z_{1})^{j}$ would be present in the PGF $V_{j+1}(z)$. For the PGF $V_{j+1}(z)$ to remain bounded, $|z_{2}|/|z_{1}|$ should be less than or equal one. Since $z_{1}=0$ and $|z_{2}|\leq|z_{1}|$, the zero $z_{2}$ must be equal to $z_{1}$, leading to pole-zero cancellation. Thus, $V_{L}(0)$ can always be determined and is a polynomial in $\Theta_{0}$ of degree $L$.
By substituting $V_{L}(0)$ in Eqn. 17, we get $W_{1}(0)$ as function of $\Theta_{0}$ of degree $(L+1)$. Further, substituting $W_{1}(0)$ in Eqn. 11 and assuming no exogenous UAS are present (that is, $E_{0}(0)=1$), we get,
$\Theta_{0}$ $=1-W_{1}(0)^{S-1}-\sum_{n=1}^{M-1}\binom{S-1}{n}(1-W_{1}(0))^{n}W_{1}(0)^{S-n-1}=h(\Theta_{0})$ (59)
The solution $\Theta_{0}$ for the equation $h(\Theta_{0})=\Theta_{0}$ is the fixed point of the polynomial function $h$. By Brouwer’s fixed-point theorem, the continuous function $h:[0,1]\rightarrow[0,1]$ has at least one fixed-point $\Theta_{0}\in[0,1]$. ∎
[…]